OpenStudy (anonymous):

Write the equation, in standard form, of the line perpendicular to 4x− 7y = 7 that passes through (2, 9)

OpenStudy (anonymous):

I got 7x-4y=-50 is this correct in standard form?

OpenStudy (anonymous):

No, that's wrong. It needs to have the same slope but it needs to cross through 2,9

OpenStudy (anonymous):

@jamone

OpenStudy (anonymous):

The answer is 7x-4y=-22

OpenStudy (anonymous):

i'm stuck how did you arrive at that answer? Thanks.

OpenStudy (anonymous):

I just graphed it. www.desmos.com/calculator

OpenStudy (anonymous):

Btw, you can give medals by clicking "Best Response".

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

y = -7x/4 + b. This line passes at point x = 2, and y = 9 9 = -7(2)/4 +b = -14/4 + b -> b = 9 +14/4 = (36 + 14)/4 = 50/4 Equation of the line: y = -7x/4 + 30/4 --> 4y + 7x = 50.

OpenStudy (anonymous):

@thu1935 thank u.